OpenStudy (anonymous):

For the function G(x) = 3^x-3 -2 , describe the asymptote. Tell how the graph is transformed from the graph of f(x) = 3^x?????

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

who is that?

OpenStudy (mathstudent55):

Is the exponent x - 3?

OpenStudy (mathstudent55):

Is the function \( \large g(x) = 3^{x - 3} - 2\)

OpenStudy (anonymous):

yes!!!

OpenStudy (mathstudent55):

Ok. Let's look at the general function \(f(x) = 3^x\)

OpenStudy (mathstudent55):

Think of x = 0, then 3^x = 3^0 = 1

OpenStudy (mathstudent55):

As the exponents become negative, for example, -1, then 3^x = 3^(-1) = 1/3 Now think of smaller and smaller negative numbers, x = -2, -3, -4, -10, -1000 3^(-2) = 1/(3^2) = 1/9 3^(-3) = 1/(3^3) = 1/27 As x becomes a smaller and smaller negative number, y gets closer and closer to zero.

OpenStudy (mathstudent55):